A vector C has magnitude of 11 m/s and makes an angle 65° degree with the positive x- axis, vector D has Cartesian co-ordinates Dx = 4m/s 2 and Dy = −3 m/s. Find the difference C -D between two vectors.


1
Expert's answer
2021-09-27T14:13:36-0400

The coordinates of "\vec{C}" :


"C_x = 11m/s\cdot \cos65\degree\\
C_y = 11m/s\cdot \sin65\degree"

The difference is given as follows:

"\vec{C} -\vec{D} = (C_x-D_x, \space C_y-D_y) = \\
=(11\cdot \cos65\degree-4,\space 11\cdot \sin65\degree+3) \approx (0.65,\space 3.9)"

Answer. "\vec{C} -\vec{D} = (0.65,\space 3.9) \space m/s".